# On the loss of context-awareness in general instruction fine-tuning

### v2

@@ 4.1.3 Benchmarks and Metrics
In addition to NIH, we report the performance on three closed-book QA tasks to benchmark context awareness: SQuAD (Rajpurkar et al., 2016), QuAC (Choi et al., 2018), and DROP (Dua et al., 2019). SQuAD is a reading comprehension benchmark where the answer to each question can be found in the context. We evaluate only the answerable subset of questions in SQuAD 1.0. QuAC is similar to SQuAD, but its questions are more open-ended and the lengths of the answers are longer. While NIH, SQuAD and QuAC only require direct retrieval from context, DROP requires a more comprehensive understanding and analysis of the given context and the answers require retrieval of multiple information from the context and discrete operations such as addition, sorting, or counting.
@@
For Needle-in-a-haystack, we report the recall error defined in Section 2.1, which is also the default metric used in Dubey et al. (2024). We set the maximum NIH context length to 1,000 for models fine-tuned on WizardLM-70K due to its shorter instruction lengths. For models fine-tuned on ShareGPT and UltraChat-200K, we set the maximum NIH context length to the maximum context window considered in fine-tuning, which is 2,000 for TinyLlama and 4,000 for Llama-2 / Llama-3. The prompt template used in NIH is the same as Section 2.1 except that we remove the response prefix and keep only the user input prompt.
@@ 4.2 Instruction Fine-tuning with Context-dependency Indicators
Dataset 0.5 0.6 0.7 0.8
ShareGPT (Vicuna) 0.14 0.10 0.07 0.04
UltraCHat-200K 0.22 0.18 0.14 0.11
WizardLM-70K 0.34 0.23 0.13 0.06
Table 4: Comparing vanilla instruction finetuning with finetuning with context-relevant indicators (+ indicator). On ‘+ Indicator’ models, [IND] is added in all evaluations. As a reference, we also list the performances evaluated on official Llama-2 and Llama-3 instruct models, which are finetuned with closed-source datasets.
SFT dataset Pretrained Model Method NIH SQuAD QuAC DROP Avg. MT-Bench ShareGPT (Vicuna) TinyLlama-1.1B Vanilla 0.9846 0.5918 0.1130 0.2739 0.4908 3.7250 + Indicator 0.9921 0.6144 0.1290 0.2784 0.5035 3.7375 Llama-2-7b Vanilla 0.3378 0.7601 0.1590 0.3390 0.3990 6.4875 + Indicator 0.7007 0.7830 0.1600 0.3390 0.4957 5.7375 Llama-3-8b Vanilla 0.8957 0.8216 0.1560 0.4215 0.5737 7.4375 + Indicator 0.9404 0.8394 0.1660 0.4317 0.5943 7.1625 UltraChat-200K TinyLlama-1.1B Vanilla 1.0 0.7289 0.1520 0.3096 0.5476 3.9000 + Indicator 1.0 0.7475 0.1570 0.3096 0.5535 4.1125 Llama-2-7b Vanilla 0.9850 0.8272 0.1540 0.3791 0.5863 5.7125 + Indicator 0.9725 0.8508 0.1570 0.3758 0.5890 5.8125 Llama-3-8b Vanilla 1.0 0.8393 0.1610 0.5099 0.6276 7.2375 + Indicator 1.0 0.8510 0.1660 0.5022 0.6298 6.8500 WizardLM-70K TinyLlama-1.1B Vanilla 0.9250 0.5994 0.0990 0.2753 0.4747 4.2750 + Indicator 0.9925 0.6279 0.1140 0.2836 0.5045 4.3000 Llama-2-7b Vanilla 0.7375 0.8229 0.1550 0.3407 0.5140 5.7750 + Indicator 0.9254 0.8260 0.1640 0.3444 0.5650 6.2250 Llama-3-8b Vanilla 0.9846 0.8765 0.1560 0.4687 0.6215 7.1125 + Indicator 0.9871 0.8792 0.1710 0.4785 0.6290 7.5250 (Closed-source datasets) Llama-2-7b-chat - 0.8264* 0.8301 0.1330 0.4422 0.5579 6.9375 Llama-3-8b-Instruct - 1.0* 0.8612 0.1850 0.4654 0.6279 8.0750 * Here NIH is evaluated without the response prefix used in Section 2.1 and the maximum context length is set to 4096 for fair comparison. Therefore, the numbers are different from the numbers in Figure 2.
@@
We compute the context-dependency scores on all three instruction fine-tuning datasets and report the ratio of identified context-dependent instructions with different threshold values $\beta$ in Table 3. Context-dependent instructions are rare in all three instruction fine-tuning datasets. This observation supports our hypothesis that the model learns an attention bias from the instruction dataset to weigh user tokens less importantly.
@@
We report the performance of NIH and three contextual QA tasks in Table 4. For the “vanilla” fine-tuning setting, we train and evaluate the model without the indicator token. For the “+ Indicator” setting, we add the indicator token to the selected subset of prompts in fine-tuning and all queries for evaluation. According to the table, “+ Indicator” outperforms “vanilla” fine-tuning in most cases. Particularly, “+ Indicator” outperforms vanilla finetuning on DROP for most cases, confirming that the training-time solution handles trade-offs between language model capabilities well. The results demonstrate that models can learn to focus more on the user-provided context when the indicator token is present in the prompt.
Conditional finetuning maintains the performance on general instruction-following.
We report the evaluation results on MT-Bench in Table 4 as an evaluation for general instruction-following. Indicators are presented during evaluation by default. Models fine-tuned with our method demonstrate comparable or even better performance on MT-Bench in most cases. Notably, MT-Bench performance increases on datasets with more context-dependent conversations (such as WizardLM-70K), which is expected when the context-dependent subset becomes more diverse. Therefore, our methods can effectively mitigate the loss of context awareness without losing the general ability to follow instructions.
@@ 5 Related Work
In this work, we define context awareness as the capability of large language models to retrieve and understand information from user-provided context and respond accordingly. Context awareness is crucial for mitigating hallucinations where the response is not consistent with the provided context, such as the “closed-domain” hallucination in Ouyang et al. (2022) and intrinsic hallucination in Huang et al. (2023). Several existing works aim to understand and mitigate these intrinsic hallucinations. Liu et al. (2024) studies the failure of context retrieval when the relevant information is in the middle of the provided context, showing the existence of positional bias in context retrieval. Follow-up work (Hsieh et al., 2024) proposes to calibrate this positional bias and mitigate the issue. In our work, we study a new role bias in popular open-source instruction-finetuned models, where the context receives less attention when marked as user tokens by the chat template.
@@ 6 Conclusion
This work highlights the detrimental effects of supervised instruction fine-tuning on the context awareness of pretrained language models, even in scenarios involving short context lengths. We have identified that the decline in context awareness is closely linked to attention allocation biases within chat templates, which are learned during conversational instruction finetuning. To address this issue, we first propose a quantitative metric to identify context-dependent instructions from the general instruction finetuning dataset. We then propose a training-time approach that employs conditional indicators to signal these instructions during the instruction-tuning process. Our method effectively maintains contextual understanding while benefiting from supervised instruction.
@@
First of all, due to computational resource limitations, our methods were only validated on smaller-sized models ($\leq$ 7B) with QLoRA (Dettmers et al., 2024). We hope to extend the experiments to full finetuning or larger models when more resources become available. Second, our technique of associating context-dependent user instructions with the indicator token may also encode other unintended styles from the selected subset of instructions. This issue is particularly aggravated when the subset of context-relevant samples is small and significantly different from the remaining dataset. One future direction is to better disentangle the context-dependency signal within the selected corpus. Finally, utilizing the indicator token requires users to know which instructions would benefit from more emphasis on the user context. Tasks demonstrated in our experiment (e.g., keyword retrieval and closed-book QA) evidently require more attention to the user context. However, complex tasks may involve multiple language skills, which complicates whether applying the indicator token will benefit the performance. A future direction is to automatically determine when to append the indicator.
@@ B.1 Full NIH Results on Open-source Official Models
Prompt templates in NIH evaluation
You are a helpful AI assistant that answers a question using only the provided document:
{context}
Question: {retrieval_question}
You are a helpful AI assistant that answers a question using only the provided context:
{context}
Question: {retrieval_question}
Document:
{context}
Answer the question accoriding to the provided document: {retrieval_question}
Context:
{context}
Answer the question accoriding to the provided context: {retrieval_question}

### v3

@@ 4.1.3 Benchmarks and Metrics
In addition to NIH, we report the performance on three closed-book QA tasks to benchmark context awareness: SQuAD (Rajpurkar et al., 2016), QuAC (Choi et al., 2018), and DROP (Dua et al., 2019). SQuAD is a reading comprehension benchmark where the answer to each question can be found in the context. We evaluate only the answerable subset of questions in SQuAD 1.0. QuAC is similar to SQuAD, but its questions are more open-ended and the lengths of the answers are longer. While NIH, SQuAD, and QuAC only require direct retrieval from context, DROP requires more complicated reasoning based on the given context, and its answers require discrete operations on the retrieved context such as addition, sorting, or counting.
@@
For Needle-in-a-haystack, we report the recall defined in Section 2.1, which is also the default metric used in Dubey et al. (2024). We set the maximum NIH context length to 1,000 for models fine-tuned on WizardLM-70K due to its shorter instruction lengths. For models fine-tuned on ShareGPT and UltraChat-200K, we set the maximum NIH context length to the maximum context window considered in fine-tuning, which is 2,000 for TinyLlama and 4,000 for Llama-2 / Llama-3. The prompt template used in NIH is the same as Section 2.1 except that we remove the response prefix and keep only the user input prompt.
@@
Table 3: Comparing vanilla instruction finetuning with finetuning with context-relevant indicators (+ indicator). On ‘+ Indicator’ models, [IND] is added in all evaluations. As a reference, we also list the performances evaluated on official Llama-2 and Llama-3 instruct models, which are finetuned with closed-source datasets. NIH, SQuAD, and QuAC are simple context-dependent tasks, while DROP and MT-Bench require more complex capabilities.
SFT dataset Pretrained Model Method NIH SQuAD QuAC DROP MT-Bench ShareGPT (Vicuna) TinyLlama-1.1B Vanilla 0.9846 59.73 15.50 27.39 3.7250 + Indicator 0.9921 62.05 17.40 27.84 3.7375 Llama-2-7b Vanilla 0.3378 76.78 23.60 33.90 6.4875 + Indicator 0.7007 79.09 24.20 33.90 5.7375 Llama-3-8b Vanilla 0.8957 83.06 24.80 42.15 7.4375 + Indicator 0.9404 84.80 24.50 43.17 7.1625 UltraChat-200K TinyLlama-1.1B Vanilla 1.0 73.03 22.70 30.96 3.9000 + Indicator 1.0 74.47 23.10 30.96 4.1125 Llama-2-7b Vanilla 0.9850 83.81 24.20 37.91 5.7125 + Indicator 0.9725 85.76 26.10 37.58 5.8125 Llama-3-8b Vanilla 1.0 85.12 25.50 50.99 7.2375 + Indicator 1.0 86.28 26.40 50.22 6.8500 WizardLM-70K TinyLlama-1.1B Vanilla 0.9250 60.51 13.80 27.53 4.2750 + Indicator 0.9925 63.39 14.60 28.36 4.3000 Llama-2-7b Vanilla 0.7375 82.89 23.70 34.07 5.7750 + Indicator 0.9254 83.13 25.30 34.44 6.2250 Llama-3-8b Vanilla 0.9846 88.25 24.60 46.87 7.1125 + Indicator 0.9871 88.53 26.00 47.85 7.5250 (Closed-source datasets) Llama-2-7b-chat - 0.8264* 83.28 22.20 44.22 6.9375 Llama-3-8b-Instruct - 1.0* 86.96 27.40 46.54 8.0750 * Here NIH is evaluated without the response prefix used in Section 2.1 and the maximum context length is set to 4096 for fair comparison. Therefore, the numbers are different from the numbers in Figure 2.
@@ 4.2 Instruction Fine-tuning with Context-dependency Indicators
Dataset 0.5 0.6 0.7 0.8 ShareGPT (Vicuna) 0.14 0.10 0.07 0.04 UltraCHat-200K 0.22 0.18 0.14 0.11 WizardLM-70K 0.34 0.23 0.13 0.06
@@
We compute the context-dependency scores on all three instruction fine-tuning datasets and report the ratio of identified context-dependent instructions with different threshold values $\beta$ in Table 4. Context-dependent instructions are rare in all three instruction fine-tuning datasets. This observation supports our hypothesis that the model learns a bias from the instruction dataset to weigh user tokens less importantly.
@@
We report the performance of NIH and three contextual QA tasks in Table 3. For the “vanilla” fine-tuning setting, we train and evaluate the model without the indicator token. For the “+ Indicator” setting, we add the indicator token to the selected subset of prompts in fine-tuning and all queries for evaluation. According to the table, “+ Indicator” outperforms “vanilla” fine-tuning in most cases. Particularly, “+ Indicator” outperforms vanilla fine-tuning consistently across different models and SFT datasets on NIH, SQuAD, and QuAC, which are more direct benchmarking tasks for context awareness performance. The results demonstrate that models can learn to focus more on the user-provided context when the indicator token is present in the prompt.
Sparsity of context-dependent instructions impacts performance on general tasks.
We report the evaluation results on MT-Bench in Table 3 as an assessment of general instruction-following capabilities. We also report results on DROP, which requires more complex reasoning capabilities beyond contextual retrieval. Indicators are presented during evaluation for “+ Indicator” models by default. Models fine-tuned with the indicator demonstrate comparable or even better performance on MT-Bench and DROP in most cases. Notably, performance on MT-Bench and DROP improves more significantly on datasets with more context-dependent conversations (such as WizardLM-70K), which maintain a more diverse distribution in their context-dependent subset.
@@ 5 Related Work
In this work, we define context awareness as the capability of large language models to retrieve and understand information from user-provided context and respond accordingly. Context awareness is crucial for mitigating hallucinations where the response is not consistent with the provided context, such as the “closed-domain” hallucination in Ouyang et al. (2022) and intrinsic hallucination in Huang et al. (2023). Several existing works aim to understand and mitigate these intrinsic hallucinations. Liu et al. (2024) studies the failure of context retrieval when the relevant information is in the middle of the provided context, showing the existence of positional bias in context retrieval. Follow-up work (Hsieh et al., 2024) proposes to calibrate this positional bias and mitigate the issue. In our work, we study a new role bias in popular open-source instruction-finetuned models, where the context receives less attention when marked as user tokens by the chat template. Some other papers such as An et al. (2024) synthesize examples targeted toward specific tasks (e.g., contextual QA) to increase the performance on context-dependent tasks. Instead, our method is more general in terms of task and input format compared. This makes our method more applicable to different types of user queries that require greater attention to user-provided context.
@@ 6 Conclusion
This work highlights the detrimental effects of supervised instruction fine-tuning on the context awareness of pretrained language models, even in scenarios involving short context lengths. We have identified that the decline in context awareness is closely linked to attention allocation biases within chat templates, which are learned during conversational instruction finetuning. To address this issue, we propose a quantitative metric to identify context-dependent instructions from the general instruction finetuning dataset. By adding indicators to the identified instructions during finetuning, we enable the model to preserve context awareness after SFT when the indicator is added during inference. Our method effectively maintains contextual understanding while benefiting from supervised instruction.
@@
First of all, due to computational resource limitations, our methods were only validated on smaller-sized models. We hope to extend the experiments to full finetuning or larger models when more resources become available. Second, our technique of associating context-dependent user instructions with the indicator token may also encode other unintended styles from the selected subset of instructions. This issue is particularly aggravated when the subset of context-relevant samples is small and significantly different from the remaining dataset. One future direction is to better disentangle the context-dependency signal within the selected corpus.
7 Impact Statement
This paper presents work aimed at advancing the field of Machine Learning. We investigate and mitigate language models’ loss of context awareness after supervised fine-tuning (SFT). Our work can potentially benefit many real-world applications, such as retrieval-augmented generation, in-context learning, and contextual question-answering. We have not identified any potential negative societal consequences.
@@ B.1 Full NIH Results on Open-source Official Models
Prompt templates in NIH evaluation You are a helpful AI assistant that answers a question using only the provided document: {context} Question: {retrieval_question} You are a helpful AI assistant that answers a question using only the provided context: {context} Question: {retrieval_question} Document: {context} Answer the question accoriding to the provided document: {retrieval_question} Context: {context} Answer the question accoriding to the provided context: {retrieval_question}